### v1

# Building Trustworthy AI: Transparent AI Systems via Large Language Models, Ontologies, and Logical Reasoning (TranspNet)

Abstract
Growing concerns over the lack of transparency in AI, particularly in high-stakes fields like healthcare and finance, drive the need for explainable and trustworthy systems. While Large Language Models (LLMs) perform exceptionally well in generating accurate outputs, their "black box" nature poses significant challenges to transparency and trust. To address this, the paper proposes the TranspNet pipeline, which integrates symbolic AI with LLMs. By leveraging domain expert knowledge, retrieval-augmented generation (RAG), and formal reasoning frameworks like Answer Set Programming (ASP), TranspNet enhances LLM outputs with structured reasoning and verification. This approach ensures that AI systems deliver not only accurate but also explainable and trustworthy results, meeting regulatory demands for transparency and accountability. TranspNet provides a comprehensive solution for developing AI systems that are reliable and interpretable, making it suitable for real-world applications where trust is critical.
1 Introduction
Symbolic AI, a fundamental branch of artificial intelligence, focuses on using structured representations of knowledge and formal logic to simulate human reasoning, problem-solving, and decision-making processes [27]. Unlike data-driven approaches such as machine learning and Large Language Models (LLMs), which derive patterns from vast amounts of unstructured data, symbolic AI models intelligence is based on rule-based systems that explicitly define rules, relationships, and logical structures. These structures include formal logic, ontologies, and semantic networks, enabling symbolic AI to work with well-defined rules to draw logical conclusions and make interpretable decisions [7].
A core component of symbolic AI is knowledge representation, where symbols denote real-world entities and their relationships. Representations in symbolic AI often employ hierarchical or graph-based structures such as semantic networks or ontologies [1, 15]. In contrast, LLMs GPT4 [22] and BERT [6], while highly effective in generating contextually relevant responses, face challenges in offering the same level of explainability due to their "black box" nature [32]. This gap is particularly concerning in high-stakes domains like healthcare, finance, and legal reasoning, where trust and transparency are paramount.
This gap between the complexity of LLMs and the demand for transparency poses a significant challenge for AI development, particularly given the legal requirements imposed by regulations like the EU’s General Data Protection Regulation (GDPR), which grants individuals the right to receive an explanation when subjected to automated decision-making processes [14].
In addition, the proposed AI Act in the European Union, for example, mandates that AI systems—particularly those relying on LLMs in high-risk applications—be designed with human oversight, transparency, and risk management at their core [10]. One of the key barriers to achieving trustworthiness with LLMs is the inherent uncertainty in their predictions, which can be influenced by factors such as noisy data, biases in training data, or insufficient training on edge cases. Addressing these uncertainties while maintaining transparency and trustworthiness is a critical challenge in developing LLM-based systems for real-world applications [23].
To address this, systems like the proposed TranspNet pipeline combine the strengths of LLMs with symbolic AI by integrating domain expert knowledge, retrieval-augmented generation (RAG), and formal reasoning frameworks like Answer Set Programming (ASP). This hybrid approach allows LLMs to benefit from the structured, logical reasoning of symbolic AI, ensuring that their outputs are not only accurate but also explainable and trustworthy [19]. Ontologies play a key role in this integration, providing a framework for verifying LLM outputs and enhancing their transparency, a characteristic central to symbolic AI [27].
By combining the power of LLMs with formal reasoning, retrieval-augmented generation, multimodal data processing, and robust documentation practices, our pipeline addresses the key challenges associated with explainability and trustworthiness in AI systems. The integration of formal logic through ASP further distinguishes our approach by providing a mechanism for verifying the logical soundness of the LLM’s outputs and addressing uncertainty in a structured and interpretable manner. Ultimately, our pipeline offers a comprehensive solution for developing LLM-based AI systems that are not only accurate but also explainable and trustworthy, meeting the needs of both academic research and industry applications.
2 State-of-the-Art in LLM Reasoning
Large language models (LLMs) have revolutionized natural language processing, achieving breakthrough performance on tasks like translation, summarization, and question-answering through in-context learning, a new paradigm that enables few-shot learning without modifying model parameters [24, 26, 29]. These models have demonstrated exceptional proficiency in what Kahneman [17] describes as “System 1” tasks – automatic, intuitive operations – but have faced challenges in “System 2” reasoning tasks, which require conscious, logical steps, such as solving math word problems [4].
Recent developments in prompt-based reasoning, such as Chain-of-Thought (CoT) prompting, have been instrumental in addressing these challenges. By guiding LLMs to generate intermediate reasoning steps, CoT has significantly improved performance on reasoning-intensive benchmarks, such as the GSM8K dataset for math word problems [29]. This approach, along with advances in self-consistency and prompt-engineering techniques, has helped bridge the gap between LLMs’ ability to perform associative tasks and their capacity for multi-step reasoning [18].
Despite these advancements, challenges remain in ensuring the faithfulness and interpretability of reasoning processes in LLMs. Techniques like self-verification and reinforcement learning have been employed to minimize errors and improve reliability, but issues such as hallucination and error accumulation persist, particularly in complex, multi-step reasoning tasks [28]. Additionally, research continues to explore how reasoning capabilities can be transferred to smaller models or embodied agents, as computational efficiency becomes an increasingly important factor in the deployment of LLMs [20].
The proposed pipeline enhances the state-of-the-art by addressing some of the core challenges faced by LLMs, particularly in ensuring explainability and trustworthiness in reasoning tasks. Unlike traditional LLM systems, the pipeline integrates domain expert knowledge, retrieval-augmented generation, and formal reasoning frameworks, enabling the system to verify and refine its outputs through external data and structured reasoning. This multi-layered approach mitigates some of the inherent issues in LLMs, such as hallucination and over-reliance on probabilistic reasoning. The inclusion of ASP [13] in the pipeline ensures that the reasoning process is grounded in logical consistency, providing verifiable and interpretable outputs. By incorporating multimodal data processing, the pipeline allows the system to handle diverse types of input, further improving its reliability and decision-making capabilities across different domains.
3 Methodology
Figure 1: Overall knowledge-driven LLM verification pipeline. The pipeline integrates domain expert knowledge, prompt engineering, retrieval-augmented generation, answer set programming, and deep learning models to ensure accurate, explainable, and trustworthy outputs from LLMs.
3.1 Overall Knowledge-Driven LLM Verification Pipeline
This pipeline integrates domain expert knowledge and structured data to enhance the reliability and LLM outputs. The process starts with expert knowledge and vocabulary analysis, followed by prompt engineering and RAG to leverage large language models for generating answers. Ontology matching ensures semantic consistency, with structured ontologies being used for mapping concepts. A critical component of the pipeline is the consciousness layer for LLMs, which incorporates ASP for applying logical reasoning. This layer enhances the verification, explanation, and trustworthiness of the generated answers, leading to a robust system aimed at providing accurate and explainable outputs [2, 9].
3.2 Domain Expert Knowledge & Data (1.a)
The pipeline begins with the input from domain experts and relevant structured data. These inputs ensure the accuracy and relevance of the information. The data may include technical documents, research papers, and domain-specific databases.
3.3 Identification of Vocabulary, Attributes, and Relationships (1.b)
The identification of vocabulary, attributes, and relationships is a crucial step in the pipeline. By defining the key terms, their attributes, and relationships, this foundational phase ensures that the LLM understands the specific context and nuances of the subject matter. This step enables accurate and meaningful interactions in subsequent stages, effectively mapping domain knowledge into a structured format to improve the comprehension and performance of the LLM.
3.4 Prompt Engineering for Consistent Triple Generation (1.c)
At this stage, the LLM is prompted to generate responses in a consistent structured format, specifically using (subject-predicate-object) triples in JSON-LD format [25]. Prompt engineering ensures that the controlled vocabulary is used consistently in the LLM output.
Prompt Example:
“Using the following vocabulary list, generate responses in the form of (subject-predicate-object) triples. Ensure all terms are used consistently according to the provided definitions: [vocabulary list].”
3.5 Retrieval-Augmented Generation (1.d)
RAG combines both retrieval-based and generation-based methods to enhance the output of LLMs. The retrieval model, typically a dense passage retriever (DPR), fetches relevant documents from a large corpus based on the input query. The generator model, usually a transformer, then uses this context to produce coherent and contextually accurate responses. This hybrid approach leverages the strengths of both retrieval and generation techniques, improving the overall relevance and quality of the generated outputs.
3.6 Large Language Model Generation (1.e)
Once the input query is refined through prompt engineering, the LLM generates structured responses in a subject-predicate-object format. The LLM’s extensive training on diverse datasets enables it to produce contextually relevant responses that maintain consistency with the controlled vocabulary defined in earlier steps. These generated triples are then mapped to the ontology, facilitating further reasoning and analysis in the pipeline.
3.7 Concept Matching (1.f)
Concept matching compares elements in the generated triples with those in the ontology to ensure alignment. Multiple matching techniques are used, including: - Name-Based Matching: String matching algorithms like Levenshtein distance and Jaccard similarity to compare element names [11]. - Structure-Based Matching: Examination of class hierarchies and property relationships in the schema [21]. - Instance-Based Matching: Comparison of actual data instances for value similarity [8]. - Linguistic Matching: Use of natural language processing techniques, such as synonym databases and word embeddings, to find semantically similar matches [31].
3.8 Ontology EL-Fragment (1.g)
Ontologies structure relationships between concepts within a domain, facilitating better data integration and analysis. The pipeline leverages EL fragments (a Description Logic EL fragment) [30], which supports ASP by enhancing the explainability and tractability of generated data. EL fragments ensure consistent terminology, efficient mapping of triples, and logical reasoning in a scalable manner, making it ideal for complex applications like biomedical informatics.
3.9 Consciousness Layer for LLM (1.h)
The role of the Consciousness Layer is to reason on the concepts extracted from (2.a) and the structured response from LLM (1.e). Therefore, the required piece of information into the domain specific ontology will be use as a knowledge-base for the ASP solver. The Consciousness Layer includes: - ASP Knowledge Base: Stores factual and procedural knowledge from mapped concepts and domain knowledge. - ASP Rules: Logical rules and constraints for reasoning over the knowledge base. - ASP Solver: The ASP solver applies logical inference to refine LLM-generated answers, ensuring consistency, accuracy, and robustness. This solver verifies that the generated triples are logically sound and contextually relevant [12].
3.10 Deep Learning Models for Feature Extraction (2.a)
To address classification, regression, clustering, and time-series problems, deep learning models could be used for feature extraction. These models are capable of extracting rich features that are then mapped to relevant concepts using techniques like DeViL (Decoding Vision features into Language) [5].
3.11 Extracted features (2.b)
Extracted features refer to the characteristics obtained after applying deep learning models for tasks such as classification, regression, and clustering. These models process data from various multimodal sources, including sensors, to extract meaningful and relevant features. This approach is particularly advantageous for our use cases, especially in the Chemical Processes context, where precise feature extraction from multimodal sensor data is crucial for accurate analysis and prediction. To further enhance the interpretability and utility of these extracted features, it is important to map them to higher-level concepts. Techniques such as the bottleneck model, inspired by the DeVil model, can be employed to identify and distill key features that contribute most significantly to the model’s performance. By identifying these critical features and relating them to specific concepts, we can achieve a more intuitive understanding of the model’s decision-making process.
4 Use-Cases
4.1 Use Case 1: Healthcare - Clinical Decision Support System
In healthcare, clinical decision support systems (CDSS) are used to assist physicians in making informed treatment decisions based on patient data and medical literature. The proposed pipeline can enhance these systems by ensuring the reliability, explainability, and trustworthiness of the LLM outputs used in patient care. The pipeline begins by integrating medical knowledge from domain experts, including data from medical research papers, clinical guidelines, patient history, and real-time data from electronic health records (EHR). This step ensures that the LLM receives accurate and up-to-date medical knowledge relevant to clinical decision-making.
In the next step, key medical terms, conditions, symptoms, diagnostic tests, and treatment options are identified and structured. This phase ensures that the LLM understands the relationships between diseases, symptoms, and treatments, facilitating more accurate clinical recommendations. Using the identified vocabulary, the LLM is prompted to generate responses related to patient conditions, diagnostics, and treatments in a structured format like (patient-symptom-disease). The system then retrieves relevant medical literature and studies using RAG to back up recommendations, ensuring that the generated outputs are grounded in evidence.
The LLM generates structured outputs, such as (patient has symptoms X, Y, Z - potential diagnosis: Disease A), based on patient data and retrieved medical information. Medical ontologies, such as SNOMED CT [3] or ICD-10 [16], are used to match the generated triples, ensuring semantic consistency and accuracy in the medical domain. The ASP-based consciousness layer ensures that the recommendations made by the LLM are logically sound and aligned with clinical guidelines, reducing the risk of incorrect recommendations. This pipeline ensures that the clinical decision support system produces trusted and explainable medical recommendations for healthcare providers, with outputs that are clinically accurate, evidence-based, and compliant with medical guidelines.
4.2 Use Case 2: Battery Design - Material Selection for Energy Storage
In battery design, selecting the right materials for components such as electrodes and electrolytes is critical to improving battery efficiency and lifespan. The proposed pipeline supports engineers and scientists by providing reliable and explainable recommendations on material combinations for energy storage systems. The pipeline starts by integrating data from scientific literature, material databases, and experimental results related to the battery technology. This input includes expert knowledge on chemical properties, performance metrics, and degradation, which guides material selection.
The next step involves identifying key attributes such as material conductivity, chemical stability, energy density, and thermal properties. Relationships between these attributes and battery performance (e.g., how conductivity affects charge/discharge rates) are mapped to ensure the LLM understands the complexity of material behavior in battery systems. The LLM is then prompted to generate suggestions for materials based on performance criteria, such as (material X - conductivity Y - potential application Z). For example, the LLM might propose materials with high conductivity and thermal stability for use in lithium-ion battery electrodes.
Next, RAG retrieves recent research papers and experimental data on materials from scientific databases. If the LLM suggests graphene as a suitable material for a battery component, RAG pulls relevant studies on graphene’s performance in energy storage applications to substantiate the recommendation. The LLM generates structured outputs suggesting material combinations in (material-property-application) triples. For instance, it might recommend using graphene for electrodes due to its high conductivity and chemical stability. Material ontologies and databases are used to match the generated triples, ensuring alignment with the known properties of materials.
Finally, the ASP-based Consciousness Layer applies logical rules to verify that the suggested materials meet the specific requirements of the battery design. For example, if a material is recommended for high-temperature environments, ASP checks whether the material’s thermal stability is sufficient, ensuring logical consistency and trustworthiness. The pipeline produces explainable, evidence-based recommendations for material selection in battery design, helping engineers and scientists optimize material choices for performance and longevity, with clear explanations of the decision-making process.
5 Conclusion
In this work, we presented a robust and comprehensive pipeline designed to enhance the explainability, accuracy, and trustworthiness of LLMs. By integrating domain expert knowledge, prompt engineering, RAG, and ASP, the proposed pipeline addresses key challenges in ensuring that LLM-generated outputs are logically consistent, contextually relevant, and semantically aligned with domain-specific knowledge.
The pipeline is structured to ensure that each stage of the process, from vocabulary identification to ontology mapping and logical reasoning, contributes to producing reliable outputs. The Consciousness Layer, with its ASP-based reasoning, represents a novel advancement in validating LLM outputs, offering a higher level of trust and transparency by providing verifiable logical consistency.
The practical application of the pipeline in diverse domains, such as healthcare and battery design, demonstrates its versatility and effectiveness in generating evidence-based, explainable, and trustworthy recommendations. In healthcare, the pipeline ensures that clinical decision support systems produce recommendations aligned with medical guidelines, ensuring patient safety and improving the quality of care. In battery design, the pipeline helps engineers make informed material selections, optimizing performance and longevity in energy storage systems.
Overall, the proposed pipeline is a significant step forward in addressing the limitations of traditional LLM outputs. It offers a comprehensive framework for generating not only accurate and context-aware answers but also ensuring that these answers are explainable, reliable, and suitable for real-world applications. As LLMs continue to evolve and are increasingly adopted across various domains, the integration of structured reasoning and expert knowledge, as demonstrated in this pipeline, will be crucial in maintaining trust and ensuring the success of AI-driven systems.